# Generalization Gap in Data Augmentation: Insights from Illumination

### v1

@@ Abstract
In the field of computer vision, data augmentation is widely used to enrich the feature complexity of training datasets with deep learning techniques. However, regarding the generalization capabilities of models, the difference in artificial features generated by data augmentation and natural visual features has not been fully revealed. This study focuses on the visual representation variable ’illumination’, by simulating its distribution degradation and examining how data augmentation techniques enhance model performance on a classification task. Our goal is to investigate the differences in generalization between models trained with augmented data and those trained under real-world illumination conditions. Results indicate that after undergoing various data augmentation methods, model performance has been significantly improved. Yet, a noticeable generalization gap still exists after utilizing various data augmentation methods, emphasizing the critical role of feature diversity in the training set for enhancing model generalization.
@@ 1 Introduction
Over the past ten years, there has been a significant revolution in computer vision field. The advancement mainly belongs to deep learning techniques, particularly the utilization of Convolutional Neural Networks (CNNs) [1, 2, 3, 4], and Transformer [5, 6] architectures. By emulating creatural visual systems, CNNs stack convolutional layers and pooling layers, while Transformers utilize self-attention mechanisms [7] to handle long-range image dependencies effectively. With all these efforts, researchers made significant progress in image classification [1, 2, 3, 4], object detection [8, 9], and semantic segmentation [10, 11]. Nowadays, Computer vision expands its applications in industry [12], healthcare [13], and transportation [14], improving convenience and quality of people’s lives.
@@
The feature diversity in visual datasets can be broadly divided into two categories: changes in the intrinsic properties of the recognition objects (e.g., the variety of vehicles in autonomous driving scenarios [22]), and changes indirectly caused by external environmental factors (e.g., different weather conditions [23] in self-driving). The diversity of features in the dataset ensures the model preserves robust recognition capabilities across various scenarios. However, when the distribution of visual features is uneven, potentially diminishes the model’s performance across various environmental conditions [24]. When data distribution imbalances occur, data augmentation is widely used as an effective method [25, 26, 27]. However, a discernible gap remains between augmented images and those captured in the real world, highlighting under certain extreme conditions, such as adverse weather, the artificial features produced by data augmentation could potentially impair model performance [28, 29]. This has prompted a reevaluation of data augmentation, whether synthetic, non-realistic pixel-wise feature characteristics might undermine a model’s generalization in real-world scenario. Our study involves controlling illumination settings in a classification task to compare model performance under real-world and data augmentation datasets, aiming reveal the effectiveness and limitations of data augmentation. Our main contributions are as follows:
  • Establishing illumination vector mapping as a data augmentation strategy for improving the generalization performance of visual models in singular distribution illumination settings;
  • By optimizing color data augmentation parameters using Bayesian optimization algorithms, we enhance the model’s generalization performance under singular illumination distribution.;
  • Through comparative experiments that focused on illumination settings as the target visual variable, we demonstrate a significant generalization gap between models trained with data augmentation datasets and those trained with real-world data.
@@ 2.1 Domain Generalization
Domain generalization aims to enhance a model’s performance on unseen data (target domain), particularly when there is a distributional discrepancy between the training data (source domain) and test data. The main strategies include domain alignment [30] to minimize distributional differences between source and target domains, meta-learning [31] which leverages learning across tasks to boost generalization capabilities, data augmentation [32, 33] that introduces sample visual diversity to enhance model robustness and ensemble learning [34] which integrates multiple models to optimize overall performance. Our study focuses on data-level augmentation as it directly enriches the visual features of training data without the need to alter network structures or training strategies, closely aligning with our goal of exploring the impact of illumination environments on model generalization.
@@ 3.1 Recognition Targets
Figure 1: An assortment of 10 distinct toy dogs serves as recognition targets in our classification task. The variety in their visual features highlights the complexity of our dataset and assesses the classification models’ ability to distinguish visual differences from subtle to pronounced.
@@
Given that this study primarily investigates the impact of data augmentation techniques on enhancing model generalization capabilities, focusing especially on the challenge of the ’illumination’ variable compared to real datasets, we aim to dive into how models adapt to visual stimuli across different data augmentation methodologies. By selecting toys with distinct visual characteristics as recognition objects, we established a foundation for subsequent research into data augmentation and model generalization.
@@ 3.3.1 Training Set
In order to thoroughly investigate the impact of the visual variable, ’illumination’, on the generalization capabilities of visual models, we purposefully constructed two training sets. The Full Spectrum Illumination Dataset (FSID) encompasses a range of illumination distributions, incorporating variations in light color and intensity. This representation ensures a uniform distribution of the ’illumination’ variable in our study. In contrast, the Singular Illumination Dataset (SID) simulates a narrowed range of illumination by selecting specific median illumination settings from FSID, transitioning from a broad, uniform distribution to a degraded, singular distribution.
@@ 3.3.2 Full Spectrum Illumination Dataset (FSID):
Figure 3: Toy 1 depicted under 15 illumination settings within the FSID, with light colors [Warm, Cool, Mixed] and intensities [-2, -1, 0, +1, +2], where illumination was measured using the light intensity meter in Fig.  2 (b), ensuring an error margin within ±20 lux.
@@ 4 Comparative Experiments
In our first experiment, we focused on how the distribution of the ’illumination’ attribute within training datasets affects the generalization performance of deep learning models. Specifically, we utilized the six deep learning models mentioned in Section 3.4 and conducted comparative experiments using the Full Spectrum Illumination Dataset (FSID) and the Singular Illumination Dataset (SID) as the training datasets. We aimed to examine the impact of differences in the distribution of the ’illumination’ variable on model generalization while keeping all other variables unchanged.
To ensure fairness and consistency for all deep learning models, we applied nearly the same training hyperparameters for all experiments. This included setting the data validation split to 0.2, the batch size to 64, choosing Adam as the optimizer, and setting the initial learning rate to 0.001. During the preprocess stage, we follow the procedural of AlexNet [1] by scaling all training images to 224*224 and making color normalization. All models were trained extensively on FSID and SID until their training loss reached convergence. Notably, CNN-based models typically converged within 10 epochs, while Transformer-based models required more training epochs. Specifically, our ViT model needed 20 epochs, and our Swin Transformer model needed 60 epochs to reach the convergence [41]. After completing all training processes, we evaluated the generalization performance of all models trained on FSID and SID using the test set described in Section  3.3.4. The evaluation metrics included model accuracy, precision, and recall on our test set, the experimental results details were shown in Tab.  1.
@@
By analyzing the results showcased in Tab.  1, when the illumination distribution degenerates to a singular one, we could clearly recognize the difference of their performance on the test set. Although all models preserved the same training configuration in FSID and SID, the four CNN and two Transformer models exhibited significant performance discrepancies due to their training sets containing different illumination distributions. The SID only contained a singular illumination distribution [Cool light, 0 level] while the FSID provided a full spectrum distribution with 15 illumination settings. Particularly, all models trained on FSID performed great performance on the test set, however, models received about 0.67 drops in accuracy when trained on SID, with similar declines observed in precision and recall. These results compellingly demonstrate that when the visual variable distribution in training sets degenerated (specifically illumination in our study), this led to a catastrophic decrease in their generalization capabilities.
@@
Figure 4: Establishing the illumination settings of FSID to generate extensive illumination vectors for augmenting the SID dataset. (a) 18% gray card, (b) scene assembled for data collection, and (c) images from the SID dataset of Toy 1, enhanced with illumination vectors under diverse illumination settings.
In this experiment, we used an 18% gray card shown in Fig.  4(a) as the subject and replicated the same 15 diverse illumination settings found in the Full Spectrum Illumination Dataset (FSID), as depicted in Fig.  4(b). Under these conditions, we captured multiple photographs. From these, we meticulously selected 100 images with consistent imaging quality for further analysis. For each selected image, we calculated the average values of the $R$, $G$, and $B$ color channels under the current lighting conditions, defining it as the environmental illumination vector $V_{\text{ill}}[R,G,B]$, where the light color $C$ includes three types (Warm, Cool, Mixed) and light intensity $I$ is divided into five levels (-2, -1, 0, +1, +2). Then, we calculated the standard illumination environment vector $V_{\text{ill,SID}}[R,G,B]$ in SID for the [Cool, 0 level] scenario and compared it with the vectors $V_{\text{ill,FSID}}^{(k)}[R,G,B]$ under the other 14 different lighting conditions. Based on these ratios, we enhanced 15,000 images in SID, selecting 100 images randomly for each lighting condition for processing. The enhancement process was accomplished by applying the corresponding illumination mapping coefficients to the global pixels of the images, thus creating the Illumination Vector Augmentation Dataset (IVAD).
@@ 4.3 Experiment 3: Color Augmentation via Bayesian Optimization in the Singular Illumination Dataset
During the training process of deep learning-based visual models, data augmentation techniques play an important role. In Experiment 2 we demonstrated that color channel enhancement could significantly improve model performance. Given that our study mainly focused on the ’illumination’ attribute, this experiment also focused on color-based data augmentation methods. We utilized Torchvision, a package from the PyTorch [43] project that provides tools and datasets for computer vision. It provided a color augmentation function ’torchvision.transforms.ColorJitter’ which included 4 variables: brightness, contrast, saturation, and hue. In this section, we aim to maximize the model’s performance on the test set by searching for the best parameter configurations on torchvision.transforms.ColorJitter data augmentation function.
To precisely adjust these four color enhancement parameters, we employed the Optuna Bayesian optimization framework [44], with the Tree-structured Parzen Estimator (TPE) [45] as the parameter sampling strategy. During the optimization process, each TPE iteration generated a new set of parameter configurations. These configurations were applied to data augmentation and subsequently used to train models according to the setup established in Experiment 1. After training, we evaluated the effectiveness of each iteration by measuring the model’s accuracy on the test set. For six different deep learning models, we conducted 200 optimization iterations for searching the best color augmentation parameter configurations. As shown in Fig.5, the results demonstrated that after 200 iterations of Bayesian optimization for data augmentation, with the best parameter configurations, the models’ performance approached or even surpassed the illumination mapping data augmentation methods used in IVAD. Our experiments showed that color enhancement techniques in Torchvision, fine-tuned through Bayesian optimization, could surpass the illumination vector mapping augmentation method in Experiment 2. However, despite all these improvements, a generalization gap still existed when compared to the FSID dataset, which included actual illumination changes. This emphasized the irreplaceable of visual feature complexity in real datasets for constructing a robust model generalization.
@@ 5 Conclusion
To conclude our study, we summarized all results in Table 3. In our research, we focused on the visual representation variable ’illumination’ by forming the Full Spectrum Illumination Dataset (FSID) with uniform distribution and the Singular Illumination Dataset (SID) with singular distribution on datasets of a classification task. In Experiment 1, we proved that when a visual representation variable degenerated to a singular distribution, it will occur a catastrophic decline in deep learning-based visual models. We performed an illumination vector mapping data augmentation method in Experiment 2. Models’ generalization abilities trained by the Illumination Vector Augmentation Dataset (IVAD) had significant improvements. Experiment 3 further conducted a Bayesian Optimization Data Augmentation (BO-DA) method, which slightly outperformed the models’ performance trained with IVAD.
Nevertheless, whether employing intuitive color mapping techniques or implementing color-based data augmentation through Bayesian optimization, these strategies still exhibit a generalization gap when compared to models trained on a complex, real-world illumination dataset. This outcome underscores the importance of covering key visual representation variables during the data collection phase and emphasizes the necessity of thoroughly considering various visual changes in experimental design and data preparation. This is crucial for achieving robust generalization capabilities in deep learning-based visual models, underscoring the crucial role of visual feature complexity in deep learning datasets.

### v2

@@ Abstract
In the field of computer vision, data augmentation is widely used to enrich the feature complexity of training datasets with deep learning techniques. However, regarding the generalization capabilities of models, the difference in artificial features generated by data augmentation and natural visual features has not been fully revealed. This study introduces the concept of "visual representation variables" to define the possible visual variations in a task as a joint distribution of these variables. We focus on the visual representation variable "illumination", by simulating its distribution degradation and examining how data augmentation techniques enhance model performance on a classification task. Our goal is to investigate the differences in generalization between models trained with augmented data and those trained under real-world illumination conditions. Results indicate that after applying various data augmentation methods, model performance has significantly improved. Yet, a noticeable generalization gap still exists after utilizing various data augmentation methods, emphasizing the critical role of feature diversity in the training set for enhancing model generalization.
@@ 1 Introduction
Over the past ten years, there has been a significant revolution in computer vision field. The advancement mainly belongs to deep learning techniques, particularly the utilization of Convolutional Neural Networks (CNNs) [1, 2, 3, 4] and Transformer [5, 6] architectures. By emulating creatural visual systems, CNNs stack convolutional layers and pooling layers, while Transformers utilize self-attention mechanisms [7] to handle long-range image dependencies effectively. With all these efforts, researchers made significant progress in image classification [1, 2, 3, 4], object detection [8, 9], and semantic segmentation [10, 11]. Nowadays, computer vision expands its applications in industry [12], healthcare [13], and transportation [14], improving convenience and quality of people’s lives.
@@
The visual diversity in datasets can be broadly divided into two categories: changes in the intrinsic properties of the recognition objects (e.g., the variety of vehicles in autonomous driving scenarios [22]), and changes indirectly caused by external environmental factors (e.g., different weather conditions [23] in self-driving). The visual diversity in the dataset ensures the model preserves robust recognition capabilities across various scenarios. However, when the distribution of visual characteristics is uneven, potentially diminishes the model’s performance across various environmental conditions [24]. When data distribution imbalances occur, data augmentation is widely used as an effective method [25, 26, 27]. However, a discernible gap remains between augmented images and those captured in the physical world, highlighting under certain extreme conditions, such as adverse weather, the artificial features produced by data augmentation could potentially impair model performance [28, 29]. This has prompted a reevaluation of data augmentation, whether synthetic, non-realistic pixel-wise feature characteristics might undermine a model’s generalization in real-world scenario. Our study involves controlling illumination settings in a classification task to compare model performance under real-world and data augmentation datasets, aiming reveal the effectiveness and limitations of data augmentation.
Figure 1: Visual representation variables decomposition guided by task prior knowledge.
In our study, we introduce the concept of "visual representation variables" to define the potential visual changes in a task as a joint distribution of these variables (Fig. 1). We focus on isolating a single visual representation variable, "illumination," to comprehensively study its effects. "Illumination" was chosen because it can be quantitatively measured and is relatively easy to replicate in data augmentation. By controlling illumination settings in a classification task, we compare model performance under real-world and data augmentation conditions to reveal the effectiveness and limitations of data augmentation.
Our main contributions are as follows:
  • We validate that the model’s generalization ability suffers a devastating impact when the illumination environment degrades into a singular distribution;
  • By using a gray card to measure scene illumination mapping and optimizing color data augmentation parameters through Bayesian optimization, we achieve significant improvements in the generalization ability of datasets with singular illumination distribution;
  • We demonstrate a significant generalization gap between models trained with data augmentation and those trained with real-world data. This emphasizes the limitations of data augmentation in replicating real-world visual features and underscores the necessity of carefully designed datasets.
@@ 2.1 Domain Generalization
Domain generalization aims to enhance a model’s performance on unseen data (target domain), particularly when there is a distributional discrepancy between the training data (source domain) and test data. The main strategies include domain alignment [30] to minimize distributional differences between source and target domains, meta-learning [31] which leverages learning across tasks to boost generalization capabilities, data augmentation [32, 33] that introduces sample visual diversity to enhance model robustness and ensemble learning [34] which integrates multiple models to optimize overall performance. The issue of domain shift in domain generalization can be decomposed into differences in the distribution of a series of "visual representation variables" between the source domains and target domains. Our study focuses on data-level augmentation as it directly enriches the visual features of training data without the need to alter network structures or training strategies, closely aligning with our goal of exploring the impact of illumination environments on model generalization.
@@ 3.1 Recognition Targets
Figure 2: An assortment of 10 distinct toy dogs serves as recognition targets in our classification task. The variety in their visual features, such as shape, color, fur texture, and attire, highlights the complexity of our dataset and assesses the classification models’ ability to distinguish visual differences from subtle to pronounced.
@@
Given that this study primarily investigates the impact of data augmentation techniques on enhancing model generalization capabilities, focusing especially on the challenge of the "illumination" variable compared to real datasets, we aim to dive into how models adapt to visual stimuli across different data augmentation methodologies. By selecting toys with distinct visual characteristics as recognition objects, we established a foundation for subsequent research into data augmentation and model generalization.
@@ 3.3.1 Training Set
In order to thoroughly investigate the impact of the visual variable, "illumination", on the generalization capabilities of visual models, we purposefully constructed two training sets. The Full Spectrum Illumination Dataset (FSID) encompasses a range of illumination distributions, incorporating variations in light color and intensity. This representation ensures a uniform distribution of the "illumination" variable in our study. In contrast, the Singular Illumination Dataset (SID) simulates a narrowed range of illumination by selecting specific median illumination settings from FSID, transitioning from a broad, uniform distribution to a degraded, singular distribution.
@@ 3.3.2 Full Spectrum Illumination Dataset (FSID):
Table 1: Illumination settings with different levels of intensity and color temperature, where illumination intensity was measured using the light intensity meter in Fig.  3 (b), ensuring an error margin within ±20 lux. And the color temperature was indirectly measured using a gray card under different illumination settings. An average value was calculated from 100 images.
Intensity -2 level -1 level 0 level +1 level +2 level Warm Light 180Lux, 3222K 540Lux, 3812K 900Lux, 4205K 1260Lux, 4388K 1620Lux, 4205K White Light 200Lux, 20397K 600Lux, 15186K 1000Lux, 12769K 1400Lux, 12527K 1800Lux, 11931K Mixed Light 400Lux, 8058K 1200Lux, 7628K 2000Lux, 7192K 2700Lux, 6607K 3500Lux, 6499K
@@
Figure 4: Toy 1 depicted under 15 illumination settings within the FSID, with light colors [Warm, Cool, Mixed] and intensities [-2, -1, 0, +1, +2]. The illumination intensity and color temperature are described in Tab.  1.
@@ 4 Comparative Experiments
In our first experiment, we focused on how the distribution of the "illumination" attribute within training datasets affects the generalization performance of deep learning models. Specifically, we utilized the six deep learning models mentioned in Section 3.4 and conducted comparative experiments using the Full Spectrum Illumination Dataset (FSID) and the Singular Illumination Dataset (SID) as the training datasets. We aimed to examine the impact of differences in the distribution of the "illumination" variable on model generalization while keeping all other variables unchanged.
To ensure fairness and consistency for all deep learning models, we applied nearly the same training hyperparameters for all experiments. This included setting the data validation split to 0.2, the batch size to 64, choosing Adam as the optimizer, and setting the initial learning rate to 0.001. During the preprocess stage, we follow the procedural of AlexNet [1] by scaling all training images to 224*224 and making color normalization. All models were trained extensively on FSID and SID until their training loss reached convergence. Notably, CNN-based models typically converged within 10 epochs, while Transformer-based models required more training epochs. Specifically, our ViT model needed 20 epochs, and our Swin Transformer model needed 60 epochs to reach the convergence [41]. After completing all training processes, we evaluated the generalization performance of all models trained on FSID and SID using the test set described in Section  3.3.4. The evaluation metrics included model accuracy, precision, and recall on our test set, the experimental results details were shown in Tab.  2.
@@
By analyzing the results showcased in Tab.  2, when the illumination distribution degenerates to a singular one, we could clearly recognize the difference of their performance on the test set. Although all models preserved the same training configuration in FSID and SID, the four CNN and two Transformer models exhibited significant performance discrepancies due to their training sets containing different illumination distributions. The SID only contained a singular illumination distribution [Cool light, 0 level] while the FSID provided a full spectrum distribution with 15 illumination settings. Particularly, all models trained on FSID performed great performance on the test set, however, models received about 0.67 drops in accuracy when trained on SID, with similar declines observed in precision and recall. These results compellingly demonstrate that when the visual variable distribution in training sets degenerated (specifically illumination in our study), this led to a catastrophic decrease in their generalization capabilities.
@@
Figure 5: Establishing the illumination settings of FSID to generate extensive illumination vectors for augmenting the SID dataset. (a) 18% gray card, (b) scene assembled for data collection, and (c) images from the SID dataset of Toy 1, enhanced with illumination vectors under diverse illumination settings (detailed in Tab. 1).
In this experiment, we used an 18% gray card shown in Fig.  5(a) as the subject and replicated the same 15 diverse illumination settings found in the Full Spectrum Illumination Dataset (FSID), as depicted in Fig.  5(b). Under these conditions, we captured multiple photographs. From these, we meticulously selected 100 images with consistent imaging quality for further analysis. For each selected image, we calculated the average values of the $R$, $G$, and $B$ color channels under the current lighting conditions, defining it as the environmental illumination vector $V_{\text{ill}}[R,G,B]$, where the light color $C$ includes three types (Warm, Cool, Mixed) and light intensity $I$ is divided into five levels (-2, -1, 0, +1, +2). Then, we calculated the standard illumination environment vector $V_{\text{ill,SID}}[R,G,B]$ in SID for the [Cool, 0 level] scenario and compared it with the vectors $V_{\text{ill,FSID}}^{(k)}[R,G,B]$ under the other 14 different lighting conditions. Based on these ratios, we enhanced 15,000 images in SID, selecting 100 images randomly for each lighting condition for processing. The enhancement process was accomplished by applying the corresponding illumination mapping coefficients to the global pixels of the images, thus creating the Illumination Vector Augmentation Dataset (IVAD).
@@ 4.3 Experiment 3: Color Augmentation via Bayesian Optimization in the Singular Illumination Dataset
During the training process of deep learning-based visual models, data augmentation techniques play an important role. In Experiment 2 we demonstrated that color channel enhancement could significantly improve model performance. Given that our study mainly focused on the "illumination" attribute, this experiment also focused on color-based data augmentation methods. We utilized Torchvision, a package from the PyTorch [43] project that provides tools and datasets for computer vision. It provided a color augmentation function "torchvision.transforms.ColorJitter" which included 4 variables: brightness, contrast, saturation, and hue. In this section, we aim to maximize the model’s performance on the test set by searching for the best parameter configurations on torchvision.transforms.ColorJitter data augmentation function.
To precisely adjust these four color enhancement parameters, we employed the Optuna Bayesian optimization framework [44], with the Tree-structured Parzen Estimator (TPE) as the parameter sampling strategy. During the optimization process, each TPE iteration generated a new set of parameter configurations. These configurations were applied to data augmentation and subsequently used to train models according to the setup established in Experiment 1. After training, we evaluated the effectiveness of each iteration by measuring the model’s accuracy on the test set. For six different deep learning models, we conducted 200 optimization iterations for searching the best color augmentation parameter configurations. As shown in Fig.5, the results demonstrated that after 200 iterations of Bayesian optimization for data augmentation, with the best parameter configurations, the models’ performance approached or even surpassed the illumination mapping data augmentation methods used in IVAD. Our experiments showed that color enhancement techniques in Torchvision, fine-tuned through Bayesian optimization, could surpass the illumination vector mapping augmentation method in Experiment 2. However, despite all these improvements, a generalization gap still existed when compared to the FSID dataset, which included actual illumination changes. This emphasized the irreplaceable of visual feature complexity in real datasets for constructing a robust model generalization.
@@ 5 Conclusion
To conclude our study, we summarized all results in Table 3. In our research, we focused on the visual representation variable "illumination" by forming the Full Spectrum Illumination Dataset (FSID) with uniform distribution and the Singular Illumination Dataset (SID) with singular distribution on datasets of a classification task. In Experiment 1, we proved that when a visual representation variable degenerated to a singular distribution, it will occur a catastrophic decline in deep learning-based visual models. We performed an illumination vector mapping data augmentation method in Experiment 2. Models’ generalization abilities trained by the Illumination Vector Augmentation Dataset (IVAD) had significant improvements. Experiment 3 further conducted a Bayesian Optimization Data Augmentation (BO-DA) method, which slightly outperformed the models’ performance trained with IVAD.
Nevertheless, whether employing intuitive color mapping techniques or implementing color-based data augmentation through Bayesian optimization, these strategies still exhibit a generalization gap when compared to models trained on a complex, real-world illumination dataset. This outcome highlights that while data augmentation can enhance model generalization, it has inherent limitations. It is essential to ensure that different visual representation variables are sufficiently complex and diverse. Most importantly, any visual representation variable should not degrade into a singular distribution. Proper dataset design is critical for achieving robust model generalization, emphasizing the importance of incorporating diverse visual features.
6 Data Limitation
One limitation of our study is the image format. Currently, images were saved in PNG format using an Intel RealSense D435i camera, which does not preserve physical illumination information as effectively as RAW format. RAW captures linear data directly from the sensor, crucial for accurate color correction and augmentation. We acknowledge that using PNG may affect our results and plan to use RAW format in future work to enhance data accuracy and robustness.